OpenStudy (anonymous):

y=-x+6 4y=x-1 solution?

OpenStudy (anonymous):

work them out simultaneously and ill let u know how ur doing :)

OpenStudy (anonymous):

could you teach me step by step please @HawkCrimson

OpenStudy (anonymous):

np :) ill show u the substitution method

OpenStudy (anonymous):

using this method, u take one equation and sub it into the other like this Since y=-x+6 then 4(-x+6)=x-1understand? i sub line 1 into line 2

OpenStudy (anonymous):

oh okay got that @HawkCrimson

OpenStudy (anonymous):

can u continue from here?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

I got -4x+24=x-1 @HawkCrimson

OpenStudy (anonymous):

perfect :D continue :))

OpenStudy (anonymous):

ok do I do variable next

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

you move it to -4x-x=-1-24, so you get -5x=-25, so x=5. Does that make sense?

OpenStudy (anonymous):

@raevenbennett understood wat anthony did?

OpenStudy (anonymous):

yes @anthony.ruhland

OpenStudy (anonymous):

To put it more succinctly, you have a system of 2 equations: y=-x+6 & 4y=x-1. If you take the 1st equation, which is solved/expressed in terms of y into the second equations y variable, you get 4(-x+6)=x-1. Then distribute the 4 in the (-x+6), which gives you -4x+24. Now solve this equation, -4x+24=x-1, for x by getting x to one side of the equal sign and the constants on the other. this yields, -5x=-25. Divide both sides by -5 and you get x=5. Now take your calculated value for x and substitute it into the 1st equation, then solve it for y. y=5+6 which yields y=11. There you have it. x=5 & y=11. This works the other way around also. Meaning you could have solved the original second equation for y, by dividing both sides by 4. This would have given you (x-1)/4. While this will work, the fractional value for y makes the math a little more work and therefore introduces a chance for mistakes over the 1st route.

OpenStudy (anonymous):

...too much words

OpenStudy (anonymous):

OK. (1)y=x+6 & 2) 4y=x-1 Take the 1st value for y and substitute in the 2nd equation's y term. 2) 4(x+6) = x-1 2) 4x + 24 = x - 1 2) -5x = -25 2) x = 5 Now plug this value for x back into the 1st equation. 1) y=x+6 1) y = 5 + 6 1) y = 11 x = 5 & y = 11

OpenStudy (anonymous):

better :)

OpenStudy (anonymous):

Thanks. I'm used to being in front of white board :-)

OpenStudy (anonymous):

these are the answers A.(3, -2) B.(-1, 5) C.(-2, 3) D.(5, 1)

OpenStudy (anonymous):

@jlightfoot

OpenStudy (anonymous):

Sorry, I was helping someone else. Where did the 4 coordinates come from? I looked back at the original question and I only saw 1 system of 2 equations. Where there more?